# Hallucinate at the Last in Long Response Generation: A Case Study on Long Document Summarization

### v2

@@ 5 How to Mitigate the Hallucinate at the Last?
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Standard Summary
Qwen2.5-7B-Instruct 1K 0.89 0.78 0.71 0.77 0.85 -6.25
2K 0.90 0.89 0.83 0.78 0.84 1.0
3K 0.87 0.84 0.79 0.83 0.82 1.25
4K 0.91 0.88 0.90 0.79 0.89 -2.0
5K 0.86 0.89 0.85 0.76 0.89 -5.0
6K 0.93 0.91 0.92 0.80 0.90 -1.0
7K 0.90 0.89 0.89 0.84 0.82 6.0
8K 0.89 0.93 0.88 0.89 0.92 -2.25
Without 1K 0.88 0.79 0.74 0.76 0.84 -4.75(+1.5)
2K 0.90 0.89 0.83 0.78 0.84 1.0(+0.0)
Sliding 3K 0.88 0.86 0.82 0.85 0.83 2.25(+1.0)
4K 0.91 0.88 0.87 0.83 0.83 4.25(+6.25)
Winodw 5K 0.87 0.87 0.81 0.82 0.80 4.25(+9.25)
6K 0.93 0.90 0.91 0.85 0.86 3.75(+4.75)
Attention 7K 0.91 0.91 0.89 0.84 0.79 9.75(+3.75)
8K 0.88 0.93 0.89 0.89 0.92 -2.25(+0.0)
Long Summary
Qwen2.5-7B-Instruct 1K 0.88 0.76 0.81 0.83 0.83 -1.0
2K 0.83 0.89 0.84 0.83 0.82 2.75
3K 0.83 0.84 0.86 0.79 0.84 -1.0
4K 0.87 0.83 0.79 0.83 0.84 -1.0
5K 0.87 0.89 0.89 0.80 0.85 1.25
6K 0.89 0.81 0.78 0.78 0.82 -0.5
7K 0.86 0.78 0.79 0.83 0.84 -2.5
8K 0.86 0.85 0.85 0.85 0.85 0.25
Without 1K 0.89 0.77 0.83 0.84 0.80 3.25(+4.25)
2K 0.83 0.89 0.83 0.80 0.78 5.75(+3.0)
Sliding 3K 0.83 0.84 0.86 0.79 0.84 -1.0(+0.0)
4K 0.87 0.83 0.79 0.83 0.84 -1.0(+0.0)
Winodw 5K 0.86 0.90 0.83 0.79 0.80 4.5(+3.25)
6K 0.89 0.82 0.84 0.83 0.78 6.5(+7.0)
Attention 7K 0.86 0.80 0.78 0.83 0.82 -0.25(+2.25)
8K 0.88 0.85 0.85 0.86 0.82 4.0(+3.75)
@@
In this experiment, we generate Wikipedia summaries with a 7K context using four methods, taking Llama3.1-8B-Instruct as the baseline. For comparison, we evaluate the following four methods: (See Appendix F for more experimental details.)
@@
  • AdaCAD Wang et al. (2025) enhances factual consistency during generation by context-aware decoding Shi et al. (2024).
@@
We report the faithfulness scores across output bins and the sensitivity results in Table 6, and the faithfulness scores across increasing output lengths for each method in Figure 7. As shown in Table 6, BooookScore achieves the lowest sensitivity (2.0), making it the closest to zero among the four methods. Moreover, as illustrated in Figure 7, BooookScore maintains the highest level of faithfulness, particularly around the 30th sentence, and notably avoids the sharp decline in faithfulness observed in other methods at the end of the summary. These results suggest that generating summaries independently for each chunk and subsequently merging them can be an effective strategy for mitigating the Hallucinate at the Last phenomenon. Nevertheless, while BooookScore serves as a practical mitigation for models without sliding window attention, it does not fundamentally resolve the issue, as it relies on chunking and merging rather than direct long-context generation. This highlights the need for future work to develop methods that maintain faithfulness when generating long text directly from long contexts.
@@ 6 Related Work
The problem of hallucination in LLMs has been a significant area of research. Prior work has largely focused on developing methods for detecting factual inconsistencies in generated text Chuang et al. (2024); Hu et al. (2024); Kim et al. (2024); Zhong and Litman (2025) and proposing strategies to mitigate their occurrence, often through improved training Zhang et al. (2023); Wan et al. (2023), decoding Shi et al. (2024); Wang et al. (2025), or prompting Zhou et al. (2023). While effective in reducing overall hallucination rates, these studies typically do not analyze the within-sequence distribution of errors.
The concept of positional bias in generation errors has been observed in related domains. Notably, studies on Large Vision-Language Models (LVLMs) have indicated that hallucinations tend to increase towards the end of generated textual descriptions of images Liu et al. (2024a); Lee et al. (2024); Min et al. (2025).
@@ 7 Conclusion
We identified and characterized the Hallucinate at the Last phenomenon in LLM-based long response generation, specifically in long document summarization. Our findings show that hallucinations disproportionately increase towards the end of long outputs, a bias amplified in longer summaries. We investigated the contributing factors and explored targeted mitigation strategies. Our work highlighted the importance of the output’s temporal dimension in LLM faithfulness and motivates future research into positionally-aware generation techniques.
@@ Acknowledgement
Appendix A Details for Generating Summaries
@@
Context Length Words Range
1K 200 to 250
2K 400 to 500
3K 600 to 750
4K 800 to 1000
5K 1000 to 1250
6K 1200 to 1500
7K 1400 to 1750
8K 1600 to 2000
Table 7: Words range used in the prompt for long output generation.
Generated Summary Length
Our experiments revealed that controlling the output length of LLMs remains a significant challenge. This difficulty becomes more pronounced as the input context length increases and as the target word range specified in the prompt becomes larger. In particular, with longer contexts, LLMs frequently generate long summaries that contain repeated sentences or even entire paragraphs. To ensure a fair and reliable evaluation, we excluded such faulty outputs and sampled only summaries with comparable lengths for use in our analysis.
Models Avg. Words
standard long
[…]
Table 8: The average word counts of both standard and long summaries generated by each model on the Wikipedia 7K dataset.
Appendix B Details for Model Usage
@@
Appendix C More Results with Different Methods
@@
Appendix D More Results on Varying Datasets
We report comparison of faithfulness scores in standard summaries generated by the Llama model at various output positions across multiple domains, with context lengths ranging from 1K to 8K in Figure 8.
@@ Appendix D More Results on Varying Datasets
Appendix E Attention Weight Calculation
$x_{full}=x_{prompt}+x_{output}=[x_{1},x_{2},...,x_{L}]$ (4)
where $L$ is the number of tokens.
$x^{k}=\{x_{i}^{k}\}_{i=1}^{100}$ (5)
for $k\in\{1,2,...,K\}$.
$x^{first}$ $=\{x^{first}_{j}\}_{j=1}^{T_{1}}$ (6)
$x^{second}$ $=\{x^{second}_{j}\}_{j=1}^{T_{2}}$
$x^{third}$ $=\{x^{third}_{j}\}_{j=1}^{T_{3}}$
where $T_{1}$, $T_{2}$, $T_{3}$ are the respective sentence lengths.
@@ Appendix F Details for Mitigation Methods
Appendix G Use of Large Language Model
This paper primarily utilized the AI assistant (ChatGPT-5) for grammar correction and refinement of clarity. It was not employed for content generation or research ideation.
@@ Appendix G Use of Large Language Model
Figure 8: Comparison of faithfulness scores in standard summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.

### v3

@@ 5 How to Mitigate the Hallucinate at the Last?
In this experiment, we generate Wikipedia summaries using four methods, taking Llama3.1-8B-Instruct as the baseline. For comparison, we evaluate the following four methods: (See Appendix F for more experimental details.)
@@
  • AdaCAD Wang et al. (2025) enhances factual consistency during generation by context-aware decoding Shi et al. (2024b).
@@
We report the faithfulness scores across output bins and the corresponding sensitivity values in Table 6, and present the bin-wise faithfulness trends for each method in Figure 7. As shown in Table 6, BooookScore achieves the lowest sensitivity (0.0), making it the closest to zero among the four methods. Moreover, as illustrated in Figure 7, BooookScore maintains the highest level of faithfulness, particularly around the final bin, and notably avoids the sharp decline in faithfulness observed in other methods at the end of the summary. These results suggest that generating summaries independently for each chunk and subsequently merging them can be an effective strategy for mitigating the Hallucinate at the Last phenomenon. In addition, we report results for other evaluation dimensions on the summaries generated using BooookScore in Table 7. The results show that BooookScore not only mitigates the Hallucinate at the Last phenomenon, but also improves the overall quality of the generated summaries.
Nevertheless, while BooookScore provides a practical mitigation for models without sliding window attention, it does not fundamentally resolve the issue, as it relies on chunking and merging rather than direct long-context generation. This underscores the need for future methods that can maintain faithfulness when generating long text directly from long contexts.
@@ 6 Related Work
The problem of hallucination in LLMs has been a significant area of research. Prior work has largely focused on developing methods for detecting factual inconsistencies in generated text Chuang et al. (2024); Hu et al. (2024); Kim et al. (2024); Zhong and Litman (2025) and proposing strategies to mitigate their occurrence, often through improved training Zhang et al. (2023); Wan et al. (2023), decoding Shi et al. (2024b); Wang et al. (2025), or prompting Zhou et al. (2023). In addition, Wan et al. (2025) analyzed positional hallucinations in long-context summarization. While these approaches reduce overall hallucination rates, most exisiting studies do not examine how hallucinations are distributed within the generated sequence, especially in long-form outputs.
Positional bias in generation errors has also been observed in related domains. In particular, studies on Large Vision–Language Models (LVLMs) report that hallucinations tend to increase toward the end of generated image descriptions Liu et al. (2024a); Lee et al. (2024); Min et al. (2025).
@@
Methods Faithfulness(%) Completeness(%) Conciseness(%)
Llama3.1-8B 90.33 100.0 41.35
BooookScore 94.33 100.0 47.83
Table 7: Overall faithfulness, completeness and conciseness scores of long summaries on the Wikipedia dataset, evaluated using FineSurE.
@@ 7 Conclusion
We identified and characterized the Hallucinate at the Last phenomenon in LLM-based long response generation, specifically in long document summarization. Our findings show that hallucinations disproportionately increase towards the end of long outputs, a bias amplified in longer summaries. We investigated the contributing factors and explored targeted mitigation strategies. Our work highlighted the importance of the output’s faithfulness and motivates future research into positionally-aware generation techniques.
@@
Ethics Statement
This study leverages publicly available datasets, including Wikipedia, Arxiv, Pubmed, Govreport, and CNNDM, to analyze long-form text generation in LLMs. All experiments were conducted in a consistent and reproducible manner across models and datasets, without any manipulation or omission of data or results. We investigate the phenomenon of Hallucinate at the Last, a tendency observed in long context-aware summarization models to generate hallucinated content toward the final portion of the summary. By drawing attention to this issue, our study contributes to ongoing efforts to enhance the reliability and factual consistency of LLM-generated summaries.
@@ Acknowledgement
Appendix A Attention Weight Calculation
$x_{full}=x_{prompt}+x_{output}=[x_{1},x_{2},...,x_{L}]$ (4)
where $L$ is the number of tokens.
$x^{k}=\{x_{i}^{k}\}_{i=1}^{100}$ (5)
for $k\in\{1,2,...,K\}$.
$x^{first}$ $=\{x^{first}_{j}\}_{j=1}^{T_{1}}$ (6)
$x^{middle}$ $=\{x^{second}_{j}\}_{j=1}^{T_{2}}$
$x^{last}$ $=\{x^{third}_{j}\}_{j=1}^{T_{3}}$
where $T_{1}$, $T_{2}$, $T_{3}$ are the respective sentence lengths.
Appendix B Details for Atomic Fact Generation
This section details our experimental details and human evaluation results from Section 2.
@@ Appendix B Details for Atomic Fact Generation
The prompt template employed for atomic fact generation is shown in Table 14.
Human Evaluation
Furthermore, we assess the quality of the atomic facts generated by the LLM through human evaluation, as reported in Table 8, focusing on atomic facts extracted from long summaries generated by Llama. For each context length, we randomly sampled 20 summaries, and the authors manually conducted the evaluation. Each atomic fact was independently annotated by the authors as either true or false, depending on whether it was supported by the corresponding summary. In total, 1,569 atomic facts were evaluated, and the two annotators agreed on 1,541 of them, corresponding to a 98.2% raw agreement. Because the task involves binary labels with a skewed class distribution, chance-corrected agreement coefficients are known to become unstable or difficult to interpret. We therefore report raw agreement as a clearer and more reliable indicator of annotation consistency in this setting.
Context Length Validation(%)
4K 95.92
5K 99.01
6K 96.18
7K 99.12
8K 99.64
Table 8: Human evaluation of the quality of atomic facts.
Appendix C Details for Model Usage
@@
Details for Bins
Throughout all experiments, we used five bins. The bins were computed using the cut function in the pandas package, and the scores of the atomic facts were assigned to bins using open interval boundaries, meaning that the boundary values were not included in the bins.
To ensure that our findings are not the result of the specific binning choice, we additionally report results using a finer-grained setting with ten bins in Table 10. The same trend persists even with this increased resolution, confirming that the Hallucinate at the Last phenomenon is robust to the choice of bin granularity.
Models Generated Summary Bins Sensitivity
1 2 3 4 5 6 7 8 9 10
Llama3.1-8B 0.80 0.79 0.83 0.77 0.71 0.74 0.72 0.83 0.76 0.70 7.22
Qwen2.5-7B 0.85 0.85 0.80 0.82 0.85 0.82 0.89 0.81 0.89 0.86 -1.78
Table 10: Faithfulness scores and sensitivity for long summaries generated by Llama and Qwen.
Appendix D More Results with Different Methods
@@
Appendix E More Results on Varying Datasets
@@ Appendix F Details for Mitigation Methods
Appendix G Details for Generating Summaries
Prompt
Context Length Words Range
4K 800 to 1000
5K 1000 to 1250
6K 1200 to 1500
7K 1400 to 1750
8K 1600 to 2000
Table 11: Words range used in the prompt for long output generation.
Generated Summary Length
Our experiments revealed that controlling the output length of LLMs remains a significant challenge. This difficulty becomes more pronounced as the input context length increases and as the target word range specified in the prompt becomes larger. In particular, with longer contexts, LLMs often produce excessively long summaries containing repeated sentences or even entire paragraphs. To ensure fair and reliable evaluation, we excluded such outputs and considered only summaries with comparable lengths in our analysis. Specifically, we compared summaries within each sentence-count range and selected those with the highest sentence counts as length-matched summaries. For example, in our setting, the length-matched summaries correspond to 28 sentences for the 4K-context case and 40 sentences for the 8K-context case.
Models Avg. Words
standard long
[…]
Table 12: The average word counts of both standard and long summaries generated by each model on the Wikipedia dataset.
Appendix H Use of Large Language Model
This paper primarily utilized the AI assistant (ChatGPT-5.1) for grammar correction and refinement of clarity. It was not employed for content generation or research ideation.
@@ Appendix H Use of Large Language Model
Input Prompt
Decompose the text into atomic facts. {{Text}}
Table 14: Prompt template used to generate atomic facts from the summary in Section 2.